OpenStudy (anonymous):

What is the estimate that results from rounding each term of 68 and 1 division symbol 7 over 12 to the nearest whole number? 68 69 35 34

OpenStudy (anonymous):

@L-Lawliet-L

OpenStudy (nurali):

68 and 1/7/12 68 and 12/7 Make it full fraction ((68*7) +12)/7 = 488/7 =69.7 rounding to the nearest whole number = 69